The sum of six times a number and 10 is 250. Find the number
Answer by harpazo(655)

Let x be the number we are looking for.
Six times a number = 6x
We now add 6x to 10.

We get 6x + 10.

The word is can be translated using the EQUAL sign.

6x + 10 = 250

6x = 250 - 10

6x = 240

x = 240/6

x = 40
The answer is 40.